OpenStudy (anonymous):

Will medal and fan best answer Billie’s monthly cell phone bill includes the cost of cell phone service, 15% tax on the cost of the cell phone service, and a $9.25 data fee. Billie’s monthly cell phone bill totals $61. What is the cost of the cell phone service before tax is added?

OpenStudy (anonymous):

42.6

OpenStudy (anonymous):

Name your variable n= cost of service your equation would be n + 0.15n + 9.25 =61 1.15n = 61-9.25 1.15n = 51.75 divide both side by 1.15 n = 45 Cost service is 45$